# Decomposition-Based Optimal Bounds for Privacy Amplification via Shuffling

### v1

@@ Abstract.
Shuffling has been shown to amplify differential privacy guarantees, offering a stronger privacy-utility trade-off. To characterize and compute this amplification, two fundamental analytical frameworks have been proposed: the privacy blanket by Balle et al. (CRYPTO 2019) and the clone paradigm (including both the standard clone and stronger clone) by Feldman et al. (FOCS 2021, SODA 2023). All these methods rely on decomposing local randomizers.
@@
In this work, we introduce a unified analysis framework—the general clone paradigm—which encompasses all possible decompositions. We identify the optimal decomposition within the general clone paradigm. Moreover, we develop a simple and efficient algorithm to compute the exact value of the optimal privacy amplification bounds via Fast Fourier Transform. Experimental results demonstrate that the computed upper bounds for privacy amplification closely approximate the lower bounds, highlighting the tightness of our approach. Finally, using our algorithm, we conduct the first systematic analysis of the joint composition of LDP protocols in the shuffle model.
Differential privacy, shuffle model, general clone, fast fourier transform, joint composition
copyright: noneccs: Security and privacy Privacy-preserving protocols
1. Introduction
@@
Differential Privacy (DP) has become a foundational framework for safeguarding individual privacy while enabling meaningful data analysis (Dwork, 2006). In real-world applications, Local Differential Privacy (LDP) is widely adopted as it eliminates the need for a trusted curator by applying noise to each user’s data before aggregation (Cormode et al., 2018; Ding et al., 2017; Yang et al., 2024a; Team, 2017). However, this decentralized approach often results in significant utility loss due to excessive noise.
@@
To address this trade-off, the shuffle model introduces a trusted shuffler between users and the aggregator (Cheu et al., 2019; Erlingsson et al., 2019; Bittau et al., 2017). The shuffler permutes the locally perturbed data, breaking the link between individual users and their submitted values. Shuffle DP retains the trust-free nature of LDP while significantly improving the privacy-utility trade-off, making it a promising model for real-world deployment (Wang et al., 2020; Imola et al., 2022; Cheu and Zhilyaev, 2022). For instance, privacy amplification by shuffling was used in Apple and Google’s Exposure Notification Privacy-preserving Analytics (Apple and Google, 2021).
@@
The amplification effect in the shuffle model means that when each of $n$ clients randomizes their data using an $\epsilon_{0}$-LDP mechanism, the shuffled reports satisfy $(\epsilon(\epsilon_{0},\delta,n),\delta)$-DP, where $\epsilon(\epsilon_{0},\delta,n)\ll\epsilon_{0}$ for sufficiently large $n$ and $\frac{1}{\delta}$ (Feldman et al., 2023a). A central theoretical challenge is to bound and compute this privacy amplification effect. A tighter bound enables the use of a larger $\epsilon_{0}$, achieving $(\epsilon,\delta)$-DP after shuffling and thereby improving utility.
@@
Among the various strategies proposed to analyze this effect (Cheu et al., 2019; Erlingsson et al., 2019; Balle et al., 2020), decomposition-based approaches have demonstrated the strongest performance (Balle et al., 2019; Feldman et al., 2022, 2023a). Two prominent frameworks are the privacy blanket by Balle et al. (Balle et al., 2019) and the clone paradigm by Feldman et al. (Feldman et al., 2022, 2023a), which includes the standard clone and the stronger clone. These approaches rely on decomposing the probability distributions induced by a local randomizer under different inputs. A decomposition naturally leads to a reduction, which yields an upper bound on the privacy amplification achieved by shuffling.
The privacy blanket provides a tailored bound for each specific local randomizer through a decomposition, and further scaling yields a looser but computable bound. It also derives a computable bound for generic randomizers in a similar way. The standard clone adopts a simple, unified decomposition, yielding a generic bound for all $\epsilon_{0}$-DP local randomizers (Feldman et al., 2022). The generic bound can be computed precisely using a numerical algorithm, and is better than the privacy blanket’s computable bound for generic randomizers.
@@
The stronger clone was proposed with a more refined decomposition, which was expected to provide a new bound for generic local randomizers and bounds for specific local randomizers (Feldman et al., 2023a). Unfortunately, a critical flaw was later found in the proof’s core lemma. A corrected version was published on arXiv (Feldman et al., 2023b), which showed that the original general bounds only hold for a restricted class of local randomizers. The original specific bounds were replaced with a version that is much weaker and cannot be computed efficiently. Although the authors conjecture that the original results may still hold in general, this remains unproven. Moreover, the flawed result has been propagated in follow-up works (Wang et al., 2024; Chen et al., 2024; Yang et al., 2024b; Wang et al., 2025).
In this paper, we provide the optimal bound for each specific randomizer among all possible decompositions, and meanwhile provide a numerical algorithm to efficiently and precisely calculate the bound.
To identify the optimal bound, we propose a unified analysis framework—the general clone paradigm-which encompasses all possible decompositions, and further show the best decomposition is the one used by the privacy blanket. However, the decomposition bound of the privacy blanket in its original form cannot be directly computed, which is why it was further scaled with Hoeffding’s and Bennett’s inequalities. Fortunately, we represent the decomposition bound in a simple form and propose a numerical algorithm using Fast Fourier Transform (FFT) to efficiently compute the bound.
With these results, we achieve the best-known bounds obtainable from decomposition-based methods. Experiments show that our computed upper bounds closely match empirical lower bounds, demonstrating the tightness of our analysis. Additionally, we discuss a potential direction to move beyond decomposition methods: identifying the exact amplification without bounding. While this approach is intuitively appealing, it currently lacks the necessary theoretical tools and remains an open problem.
@@
Finally, we conduct the first systematic analysis of joint composition in the shuffle model using our algorithm. In classical DP, $k$-fold composition refers to applying $k$ independent mechanisms to the same dataset:
@@
In contrast, our notion of joint composition applies independent mechanisms to different datasets:
@@
Joint composition is widely used in LDP applications such as joint distribution estimation and heavy hitter detection (Domingo-Ferrer and Soria-Comas, 2022; Ren et al., 2018; Bassily and Smith, 2015; Kikuchi, 2022). For example, when each user’s data contains $d$ attributes, applying an $\frac{\epsilon}{d}$-LDP mechanism to each attribute ensures overall $\epsilon$-LDP while preserving inter-attribute correlations.
While existing studies have analyzed $k$-fold composition in the shuffle model (Koskela et al., 2021, 2020), we focus on the case where the local mechanism itself is a joint composition. Our experimental results show that existing methods yield relatively loose bounds in this setting, whereas our algorithm computes significantly tighter results by leveraging the optimal bounds.
Our contributions can be summarized as follows:
  • We propose the general clone paradigm, which subsumes all decomposition-based methods, and identify the optimal bound it can provide for a specific local randomizer.
  • We provide an efficient numerical algorithm for computing the optimal bounds via FFT.
  • We perform the first systematic analysis of joint composition in the shuffle model, achieving significantly tighter amplification bounds.
2. Preliminaries
@@
Differential privacy is a privacy-preserving framework for randomized algorithms. Intuitively, an algorithm is differentially private if the output distribution does not change significantly when a single individual’s data is modified. This ensures that the output does not reveal substantial information about any individual in the dataset. The hockey-stick divergence is commonly used to define $(\epsilon,\delta)$-DP.
@@
We say that $P$ and $Q$ are $(\epsilon,\delta)$-indistinguishable if:
$\max\{D_{e^{\epsilon}}(P\parallel Q),D_{e^{\epsilon}}(Q\parallel P)\}\leq\delta.$
@@
An algorithm $\mathcal{R}:\mathbb{X}^{n}\to\mathbb{Z}$ satisfies $(\epsilon,\delta)$-differential privacy if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, $\mathcal{R}(X)$ and $\mathcal{R}(X^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
@@
An algorithm $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ satisfies local $(\epsilon,\delta)$-differential privacy if for all $x,x^{\prime}\in\mathbb{X}$, $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
Here, $\epsilon$ is referred to as the privacy budget, which controls the privacy loss, while $\delta$ allows for a small probability of failure. When $\delta=0$, the mechanism is also called $\epsilon$-DP.
Following conventions in the shuffle model based on randomize-then-shuffle (Balle et al., 2019; Cheu et al., 2019), we define a single-message protocol $\mathcal{P}$ in the shuffle model as a pair of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{O}$. We call $\mathcal{R}$ the local randomizer, $\mathbb{Y}$ the message space of the protocol, $\mathcal{A}$ the analyzer, and $\mathbb{O}$ the output space.
@@
A protocol $\mathcal{P}=(\mathcal{R},\mathcal{A})$ satisfies $(\epsilon,\delta)$-differential privacy in the shuffle model if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, the distributions $\mathcal{S}\circ\mathcal{R}^{n}(X)$ and $\mathcal{S}\circ\mathcal{R}^{n}(X^{\prime})$ are $(\epsilon,\delta)$-indistinguishable.
3. Review of Existing Analysis Techniques
In this section, we review existing analysis techniques for studying privacy amplification in the shuffle model. We begin by introducing the standard clone paradigm due to its simplicity (Feldman et al., 2022). Next, we restate the privacy blanket framework, using consistent terminology with the former (Balle et al., 2019). On one hand, this clarifies misconceptions about the privacy blanket framework found in subsequent literature (Koskela et al., 2021). On the other hand, it aids in identifying the intrinsic connection between the two approaches, which will be explored in Section 4.
We then discuss the subsequent attempts to extend the clone paradigm, specifically the stronger clone. We show the vision and failure of both the original and corrected versions of the stronger clone.
3.1. Standard clone
The intuition behind the standard clone paradigm is as follows (Feldman et al., 2022): Suppose that $X^{0}$ and $X^{1}$ are neighbouring databases that differ on the first datapoint, $x_{1}^{0}\neq x_{1}^{1}$. A key observation is that for any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}$ and data point $x$, $\mathcal{R}(x)$ can be seen as sampling from the same distribution as $\mathcal{R}(x_{1}^{0})$ with probability at least $e^{-\varepsilon_{0}}/2$ and sampling from the same distribution as $\mathcal{R}(x_{1}^{1})$ with probability at least $e^{-\varepsilon_{0}}/2$. That is, with probability $e^{-\varepsilon_{0}}$ each data point can create a clone of the output of $\mathcal{R}(x_{1}^{0})$ or a clone of $\mathcal{R}(x_{1}^{1})$ with equal probability. Thus $n-1$ data elements effectively produce a random number of clones of both $x_{1}^{0}$ and $x_{1}^{1}$, making it more challenging to distinguish whether the original dataset contains $x_{1}^{0}$ or $x_{1}^{1}$ as its first element.
Due to the $\epsilon_{0}$-DP property of the local randomizer $\mathcal{R}$, we have the following inequality:
$\forall x_{i}\in\mathbb{X},\forall y\in\mathbb{Y}:\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\epsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{0})=y]$
$\wedge\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\epsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{1})=y].$

### v5

@@ Abstract
Shuffling has been shown to amplify differential privacy guarantees, enabling a more favorable privacy-utility trade-off. To characterize and compute this amplification, two fundamental analytical frameworks have been proposed: the privacy blanket by Balle et al. (CRYPTO 2019) and the clone—including both the standard and stronger variants—by Feldman et al. (FOCS 2021, SODA 2023). These frameworks share a common foundation: decomposing local randomizers into structured components for analysis.
@@
In this work, we introduce a unified analytical framework—the general clone paradigm—which subsumes all possible decompositions, with the clone and blanket decompositions arising as special cases. Within this framework, we identify the optimal decomposition, which is precisely the one used by the privacy blanket. Moreover, we develop a simple and efficient algorithm based on the Fast Fourier Transform (FFT) to compute optimal privacy amplification bounds. Experimental results show that our computed upper bounds nearly match the lower bounds, demonstrating the tightness of our method. Building on this method, we also derive optimal amplification bounds for both joint and parallel compositions of LDP mechanisms in the shuffle model.
1 Introduction
@@
Differential Privacy (DP) has become a foundational framework for safeguarding individual privacy while enabling meaningful data analysis [Dwo06]. In real-world applications, Local Differential Privacy (LDP) is widely adopted as it eliminates the need for a trusted curator by applying noise to each user’s data before aggregation [CJK+18, DKY17, YGZ+24, DPT17]. However, this decentralized approach often results in significant utility loss due to excessive noise.
@@
To address this trade-off, the shuffle model introduces a trusted shuffler between users and the aggregator [CSU+19, EFM+19, BEM+17]. This work considers the single-message shuffle model, where each client sends a report of their data and these reports are then anonymized and randomly shuffled before being sent to the server [FMT23a]. Shuffle DP improves the privacy-utility trade-off while minimizing trust assumptions to a single shuffler, making it a promising model for real-world deployment [WDX+20, IMC22, CZ21]. For instance, privacy amplification by shuffling was used in Apple and Google’s Exposure Notification Privacy-preserving Analytics [AG21].
@@
The “amplification-by-shuffling” theorem in the shuffle model implies that when each of the $n$ users randomizes their data using an $\varepsilon_{0}$-LDP mechanism, the collection of shuffled reports satisfies $(\varepsilon(\varepsilon_{0},\delta,n),\delta)$-DP, where $\varepsilon(\varepsilon_{0},\delta,n)\ll\varepsilon_{0}$ for sufficiently large $n$ and not too small $\delta$ [FMT23a]. A central theoretical challenge is to characterize and compute this privacy amplification effect. A tighter bound allows for a larger value of $\varepsilon_{0}$ while still achieving $(\varepsilon,\delta)$-DP after shuffling, thereby improving the utility of the mechanism. It also enables a more accurate analysis of the overall $(\varepsilon,\delta)$-guarantees resulting from shuffling $n$ independent $\varepsilon_{0}$-LDP reports.
Researchers are interested in two types of upper bounds on privacy amplification. The first is a generic bound, which provides a uniform guarantee for all $\varepsilon_{0}$-DP local randomizers. The second is a specific bound, which is tailored for a given $\varepsilon_{0}$-DP local randomizer. Generic bounds are primarily studied from a theoretical perspective, as they allow the derivation of asymptotic expressions for $\varepsilon(\varepsilon_{0},n,\delta)$ [EFM+19, BBGN19, FMT22]. In contrast, specific bounds are usually much tighter and are of significant practical importance, as they provide precise guarantees for concrete mechanisms used in real-world applications.
@@
Among the various strategies proposed to analyze this effect [CSU+19, EFM+19, BKM+20], decomposition-based approaches are the most popular [BBGN19, FMT22, FMT23a]. Two prominent frameworks are the privacy blanket by Balle et al. [BBGN19] and the clone paradigm by Feldman et al., which includes the standard clone [FMT22] and the stronger clone [FMT23a]. These approaches rely on decomposing the probability distributions induced by a local randomizer under different inputs. A decomposition naturally leads to a reduction, which yields an upper bound on the privacy amplification via shuffling.
The privacy blanket framework designs a tailored decomposition to obtain a bound for each specific local randomizer [BBGN19]. However, this bound cannot be computed precisely in its original form. To address this, the framework resorts to further approximations—such as Hoeffding’s and Bennett’s inequalities—which yield looser but computable bounds. A computable generic bound is also derived in a similar manner.
The standard clone adopts a simple, unified decomposition that yields a generic bound applicable to all $\varepsilon_{0}$-DP local randomizers [FMT22]. This bound can be computed precisely via a dedicated numerical algorithm, and empirically outperforms the computable generic bound derived from the privacy blanket framework. However, the standard clone does not provide specific bounds, except in the special case of $k$-ary Randomized Response.
@@
The stronger clone was introduced with a more refined decomposition and was expected to yield improved bounds for both generic and specific local randomizers [FMT23a]. Unfortunately, a critical flaw was later identified in the proof’s core lemma. A corrected version was released on arXiv [FMT23b], which showed that the original generic bounds hold only for a restricted class of local randomizers. The specific bounds were also revised and replaced with a version that is much weaker and lacks an efficient computation method.
1.1 Our contributions
In this paper, we focus on computing specific bounds. We provide the optimal bound achievable via any decomposition-based method for a given randomizer. Furthermore, we introduce an efficient numerical algorithm that can compute this optimal bound. In addition, we present methods for computing optimal amplification bounds for both joint composition and parallel composition in the shuffle model, achieving substantially tighter results than existing approaches.
General Clone and Optimal Bound.
In prior work on Shuffle DP, the specific bounds of the blanket framework has seen limited use; instead, its associated analysis has mostly been applied in specific cases when the blanket distribution is uniform, as in $k$-ary randomized response ($k$-RR) [BBGN19], optimal local hash [WDX+20] or in multi-message shuffle models where uniform blanket noise is directly injected [LWY22]. Although the clone and blanket appear somewhat similar, no formal connection between them has been established.
In this work, we present a unified analytical framework—termed the general clone—which encompasses all possible decompositions, including both the clone and the blanket as special cases. We further demonstrate that, among all such decompositions, the optimal one corresponds precisely to the privacy blanket.
Efficient Computation of Optimal Bounds.
The decomposition bound derived from the privacy blanket, in its original form, is not directly computable. To address this, we reinterpret it through the lens of the general clone framework and derive a simplified representation that admits efficient numerical evaluation via the Fast Fourier Transform (FFT).
Specifically, we extend the Privacy Amplification Random Variable (PARV) introduced in the privacy blanket literature to define the Generalized Privacy Amplification Random Variable (GPARV). In this formulation, the privacy loss corresponds to the expected positive part of the sum of $n$ independent and identically distributed GPARVs $G_{i}$:
$\frac{1}{n}\mathbb{E}\left[\max\left\{0,\sum_{i=1}^{n}G_{i}\right\}\right].$
This expression naturally takes the form of a convolution, enabling efficient evaluation via FFT.
Our method yields the best-known bounds achievable through decomposition-based techniques. Empirical results show that the computed upper bounds closely match the corresponding lower bounds, demonstrating both tightness and reliability. Notably, the new bounds improve the resulting $\varepsilon$ by at least 10% across all settings, and by up to 50% when $\varepsilon_{0}$ is large and $n$ is small.
Optimal bounds for joint composition.
@@ 1.1 Our contributions
We conduct the first systematic analysis of joint composition in the shuffle model using our algorithm. In classical DP, $k$-fold composition refers to applying $k$ independent mechanisms to the same data:
@@
In contrast, our notion of joint composition applies independent mechanisms to different data:
@@
These two definitions are interconvertible: treating $(D_{1},D_{2},\ldots,D_{k})$ as a single data $D$ yields joint composition as a special case of $k$-fold composition; conversely, setting $D_{1}=D_{2}=\cdots=D_{k}=D$ makes $k$-fold composition a special case of joint composition. Our analysis of joint composition also applies to the $k$-fold composition setting. However, the joint composition framework captures a more general scenario—namely, when $D_{1},D_{2},\dots,D_{k}$ are not necessarily identical and may even belong to different domains. It is important to note that our setting concerns the composition of LDP mechanisms in the single-message shuffle model, in which $D$ or the tuple $(D_{1},D_{2},\dots,D_{k})$ corresponds to the data held by an individual user. Moreover, during the shuffling process, these outputs are treated as a single tuple and are shuffled together as a unit.
@@
Joint composition is widely used in LDP applications such as joint distribution estimation and heavy hitter detection [DFSC22, RYY+18, BS15, Kik22]. For example, when each user’s data contains $d$ attributes, applying an $\frac{\varepsilon_{0}}{d}$-LDP mechanism to each attribute ensures overall $\varepsilon_{0}$-LDP while preserving inter-attribute correlations. Our experimental results show that existing methods yield relatively loose bounds in this setting, whereas our algorithm computes significantly tighter results by leveraging the optimal bounds.
Optimal bounds for parallel composition.
Additionally, we show how to compute the optimal bounds for the parallel composition of LDP mechanisms. Parallel composition refers to the setting where each user randomly selects a differentially private mechanism $\mathcal{R}_{i}$ with probability $p_{i}$ and applies it to their own data:
$\mathcal{R}_{[1,m]}^{\parallel,\boldsymbol{p}}(x)=(i,\mathcal{R}_{i}(x)),w.p.\ p_{i},\quad i=1,2,\dots,m$
where $\boldsymbol{p}=(p_{1},\dots,p_{m})$ is a probability vector with $\sum_{i=1}^{m}p_{i}=1$.
It is particularly suited for scenarios in which the analyst aims to compute multiple statistical queries—each corresponding to a distinct local mechanism $\mathcal{R}_{1},\dots,\mathcal{R}_{m}$. Parallel composition partitions the user population into $m$ non-overlapping subsets, assigning each subset to one of the $m$ estimation tasks. As a result, each subset can utilize the full privacy budget $\varepsilon_{0}$, rather than splitting the budget across all tasks. This technique has been widely adopted in the literature [BNST17, WXY+19, WDZ+19, WLJ21, CKS18] to achieve better utility compared to naive budget division.
It is worth noting that parallel composition also characterizes Poisson subsampling in the shuffle model [BBG18, Ste22, BBG20]: each user participates in a mechanism $\mathcal{R}^{*}$ with probability $p^{*}$, which is equivalent to the parallel composition $\mathcal{R}_{[1,2]}^{\parallel,\boldsymbol{p}}(x)$, where $\mathcal{R}_{1}(x)=\mathcal{R}^{*}(x)$, $\mathcal{R}_{2}(x)=\bot$, and $p_{1}=p^{*}$, $p_{2}=1-p^{*}$.
To compute the privacy amplification of parallel composition in the shuffle model was previously studied in [WPL+24], but their analysis inherited a technical flaw from [FMT23a], resulting in incorrect conclusions. Our method provides the first correct computation of the optimal bound in this setting.
1.2 Related Work
We now discuss prior work most relevant to the technical contributions of our paper. Koskela et al. [KHH23] investigate $k$-fold composition of Shuffle DP protocols. Their approach computes the Privacy Loss Random Variable (PLRV) of a dominating pair for a single-round shuffle protocol, and then applies the FFT to analyze multiple rounds (i.e., composition over $k$ executions) [KJH20]. In contrast, our work focuses on joint composition, where a single-round shuffle is applied to a composite local randomizer constructed from multiple sub-randomizers.
Some subtle aspects warrant further clarification, particularly regarding how coupling among sub-randomizers affects the composition analysis. In the single-message shuffle model, when the local randomizer is a joint composition, the outputs of the sub-local randomizers for a single user are inherently coupled. From a theoretical standpoint, privacy analysis in this setting ($\mathcal{S}\circ(\mathcal{M}_{1}\times\mathcal{M}_{2}\times\dots\times\mathcal{M}_{k})^{n}$) is not equivalent to applying composition to the individual mechanisms $\mathcal{R}_{i}$ after shuffling $(\mathcal{S}_{1}\circ\mathcal{M}_{1}^{n},\mathcal{S}_{2}\circ\mathcal{M}_{2}^{n},\dots,\mathcal{S}_{k}\circ\mathcal{M}_{k}^{n})$. Furthermore, the standard compositional properties of the PLRV do not apply.
In [KHH23], computing the PLRV requires a dominating pair [ZDW22], obtained either from the stronger clone (in the general case) or from the privacy blanket (in the case of $k$-RR). In the general case, [KHH23] inherits a flaw that originates from the original version of the stronger clone construction (see Section 3.3). For any specific local randomizer, our work demonstrates that the privacy bound derived from the blanket framework is strictly tighter than that obtained via the clone.
Theoretically, computing the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ (as defined in Section 3.2) and applying the Hockey-Stick divergence $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})$ yields the same amplification bound as computing it directly from the GPARV; both approaches evaluate $D_{e^{\varepsilon}}(P_{0}^{\mathsf{B}}\parallel P_{1}^{\mathsf{B}})$. However, for general local randomizers, the PLRV of $(P_{0}^{\mathsf{B}},P_{1}^{\mathsf{B}})$ is analytically intractable due to the complex structure of the blanket distribution. GPARV overcomes this limitation by expressing the privacy loss in a convolution-friendly form, enabling efficient FFT-based evaluation even in the absence of a closed-form expression for the PLRV.
1.3 Roadmap
Section 2 introduces the definitions of differential privacy and the shuffle model. In Section 3, we briefly review the clone and blanket frameworks. Section 4 presents our general clone analysis framework, demonstrating that both the clone and blanket decompositions are special cases within it. We prove the optimality of the blanket decomposition and, due to the redundancy of the original construction, propose a method for computing the Simplified Optimal Decomposition. For both joint and parallel compositions, we show that their optimal decompositions are given by the Cartesian product and the weighted combination of the optimal decompositions of their respective sub-local randomizers.
Section 5 introduces the Generalized Privacy Amplification Random Variable (GPARV) and an FFT-based algorithm for computing the optimal privacy amplification bound. We also clarify the relationship between GPARV and the original PARV introduced by Balle et al. [BBGN19]. Section 6 presents our experimental results.
Sections 7 and 8 discuss potential directions beyond decomposition-based techniques, as well as limitations of the current work. Finally, Appendix A provides detailed derivations for computing the optimal decompositions of several commonly used local randomizers.
2 Preliminaries
@@
Differential privacy is a privacy-preserving framework for randomized algorithms. Intuitively, an algorithm is differentially private if the output distribution does not change significantly when a single individual’s data is modified. This ensures that the output does not reveal substantial information about any individual in the dataset. The hockey-stick divergence is commonly used to define $(\varepsilon,\delta)$-DP.
@@
We say that $P$ and $Q$ are $(\varepsilon,\delta)$-indistinguishable if:
$\max\{D_{e^{\varepsilon}}(P\parallel Q),D_{e^{\varepsilon}}(Q\parallel P)\}\leq\delta.$
@@
An algorithm $\mathcal{R}:\mathbb{X}^{n}\to\mathbb{Z}$ satisfies $(\varepsilon,\delta)$-differential privacy if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, $\mathcal{R}(X)$ and $\mathcal{R}(X^{\prime})$ are $(\varepsilon,\delta)$-indistinguishable.
@@
An algorithm $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ satisfies local $(\varepsilon,\delta)$-differential privacy if for all $x,x^{\prime}\in\mathbb{X}$, $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ are $(\varepsilon,\delta)$-indistinguishable.
Here, $\varepsilon$ is referred to as the privacy budget, which controls the privacy loss, while $\delta$ allows for a small probability of failure. When $\delta=0$, the mechanism is also called $\varepsilon$-DP.
Following conventions in the shuffle model based on randomize-then-shuffle [BBGN19, CSU+19], we define a single-message protocol $\mathcal{P}$ in the shuffle model as a pair of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{O}$. We call $\mathcal{R}$ the local randomizer, $\mathbb{Y}$ the message space of the protocol, $\mathcal{A}$ the analyzer, and $\mathbb{O}$ the output space.
@@
A protocol $\mathcal{P}=(\mathcal{R},\mathcal{A})$ satisfies $(\varepsilon,\delta)$-differential privacy in the shuffle model if for all neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, the distributions $\mathcal{S}\circ\mathcal{R}^{n}(X)$ and $\mathcal{S}\circ\mathcal{R}^{n}(X^{\prime})$ are $(\varepsilon,\delta)$-indistinguishable.
3 Review of Existing Analysis Techniques
In this section, we review existing analysis techniques for studying privacy amplification in the shuffle model. We begin by introducing the standard clone due to its simplicity [FMT22]. Next, we restate the privacy blanket framework, using consistent terminology with the former [BBGN19]. It aids in identifying the intrinsic connection between the two approaches, which will be explored in Section 4.
We then discuss subsequent attempts to extend the clone paradigm, specifically the stronger clone. We highlight both the vision and the shortcomings of its original and corrected versions. The lessons learned from these failures are incorporated into the design of the general clone framework (see Section 4.1).
3.1 Standard clone
The intuition behind the standard clone paradigm is as follows [FMT22]: Suppose that $X^{0}$ and $X^{1}$ are neighbouring databases that differ on the first datapoint, $x_{1}^{0}\neq x_{1}^{1}$. A key observation is that for any $\varepsilon_{0}$-DP local randomizer $\mathcal{R}$ and data point $x$, $\mathcal{R}(x)$ can be seen as sampling from the same distribution as $\mathcal{R}(x_{1}^{0})$ with probability at least $e^{-\varepsilon_{0}}/2$ and sampling from the same distribution as $\mathcal{R}(x_{1}^{1})$ with probability at least $e^{-\varepsilon_{0}}/2$. That is, with probability $e^{-\varepsilon_{0}}$ each data point can create a clone of the output of $\mathcal{R}(x_{1}^{0})$ or a clone of $\mathcal{R}(x_{1}^{1})$ with equal probability. Thus $n-1$ data elements effectively produce a random number of clones of both $x_{1}^{0}$ and $x_{1}^{1}$, making it more challenging to distinguish whether the original dataset contains $x_{1}^{0}$ or $x_{1}^{1}$ as its first element.
Due to the $\varepsilon_{0}$-DP property of the local randomizer $\mathcal{R}$, we have the following inequality:
$\forall x_{i}\in\mathbb{X},\forall y\in\mathbb{Y}:\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\varepsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{0})=y]$
$\wedge\Pr[\mathcal{R}(x_{i})=y]\geq\frac{1}{e^{\varepsilon_{0}}}\Pr[\mathcal{R}(x_{1}^{1})=y].$